What is the pH of a buffer solution prepared by dissolving 0.40 mols HC3H3O2 and 0.35...

  1. What is the pH of a buffer solution prepared by dissolving 0.40 mols HC3H3O2 and 0.35 mols C3H3O2- in 250 mL water? ( pKa (HC3H3O2) = 4.25 )
Answer #1

Moles of HC3H3O2 = 0.40 mol

and Moles C3H3O2- =0.35 mol  

pKa (HC3H3O2) = 4.25

We know Henderson equation for buffer is pH = pKa + log [C3H3O2- / HC3H3O2]

                                                                   = 4.25 + log [0.35 / 0.4 ]

                                                                    = 4.25 + log (0.875) = 4.25 + (-0.05799)= 4.19